OpenStudy (4meisu):

The mean and standard deviation of a group of 20 boys height's are 165 cm and 6.1 cm respectively. If a boy who is 1.76m leaves the group find the new mean and variance of a sample of 62 people.

OpenStudy (campbell_st):

well you set up an equation Mean = sum of the score/number of scores so 165 = sum of scores/ 20 so the total for the sum of the scores is 165 * 20 then is 1.76 tall person leaves... subtract 176 from the total so the new mean will be Mean = (165 * 20 - 176)(19) I hope this makes sense...

OpenStudy (campbell_st):

the sum of the deviations will be change as (176- 165) is removed from the data... so using the standard deviation fromula \[SD = \sqrt{\frac{1}{n} \sum_{n}^{i = 1}(x_{i} - \mu)^2}\] so \[6.1 = \sqrt{\frac{1}{20}\sum_{x_{i} = 1}^{20}(x_{i} - 165)^2}\] then \[6.1 ^2 \times 20 = \sum_{i = 1}^{20} (x_{i} - 165)^2\] so \[\sum_{i=1}^{20} (x_{i} - 165)^2 = 744.2\] so you need to subtract (176 - 165)^2 to find the new total of the deviations squared. so the sum of the deviations is 623.2 then \[SD = \sqrt{\frac{1}{19} \times 623.2}\] hope this helps... you just need to finish the calculations.